OpenStudy (anonymous):

Solve the equation sin^2x - sqroot2 cosx = cos^2x + sqroot2 cosx + 2 for x in the interval 0 to 2pi

OpenStudy (campbell_st):

well use the trig identity \[\sin^2(x) + \cos^2(x) = 1\] make sin^2(x) the subject and then substitute into the left hand side of the equation.. then you'll end up with a quadratic equation after collecting like terms.. just solve the quadratic

OpenStudy (anonymous):

So would it simplify to \[2\cos^2(x) + 2\sqrt{2} \cos(x)+1=0\]?

OpenStudy (campbell_st):

yep....thats it... I'd use the general quadratic formula now to find cos(x) and then you can find x

OpenStudy (anonymous):

but in the quadratic equation it becomes \[\frac{-2\sqrt{2}\pm \sqrt{0}}{4}\]

OpenStudy (campbell_st):

yep so \[\cos(x) = \frac{-\sqrt{2}}{2}\] which is think is 2nd and 3rd quadrants.... 3/4 pi and 5/4 pi

OpenStudy (anonymous):

Ohhh that makes sense. Thank you so much! :) Putting it into the calculator confused me.

OpenStudy (campbell_st):

lol... it can be like that... well done...